# MonoMobility: Zero-Shot 3D Mobility Analysis from Monocular Videos

### v2

@@ Abstract
Accurately analyzing the motion parts and their motion attributes in dynamic environments is crucial for advancing key areas such as embodied intelligence. Addressing the limitations of existing methods that rely on dense multi-view images or detailed part-level annotations, we propose an innovative framework that can analyze 3D mobility from monocular videos in a zero-shot manner. This framework can precisely parse motion parts and motion attributes only using a monocular video, completely eliminating the need for annotated training data. Specifically, our method first constructs the scene geometry and roughly analyzes the motion parts and their initial motion attributes combining depth estimation, optical flow analysis and point cloud registration method, then employs 2D Gaussian splatting for scene representation. Building on this, we introduce an end-to-end dynamic scene optimization algorithm specifically designed for articulated objects, refining the initial analysis results to ensure the system can handle ‘rotation’, ‘translation’, and even complex movements (‘rotation+translation’), demonstrating high flexibility and versatility. To validate the robustness and wide applicability of our method, we created a comprehensive dataset comprising both simulated and real-world scenarios. Experimental results show that our framework can effectively analyze articulated object motions in an annotation-free manner, showcasing its significant potential in future embodied intelligence applications.
@@ 2 Related Work
3D Gaussian-Based Dynamic Scene Reconstruction. 3D Gaussian-based approaches [44, 35, 40] achieve state-of-the-art performance in 3D/4D reconstruction. 4D Gaussian Splatting [44] predicts Gaussian attribute changes over time using a six-plane dynamic module. Subsequent works [35, 40] explore single-view Gaussian dynamic reconstruction. DGMarbles [35] and Shape of Motion [40] leverage advanced data-driven techniques to obtain geometric and motion priors. Unlike these works, which treat each Gaussian independently, our method unifies motion part Gaussians to capture realistic motion patterns.
@@ 4.3 Initialization
In Section 4.2, we have obtained the point cloud sequence $\{P_{i}^{k}\}_{i=1}^{N}$ for each motion part Gaussians $GS^{k}$. However, due to errors in geometric and spatiotemporal alignment of depth estimation results, only when the movement is significant, the motion pattern can be roughly reflected in the transformation between two frame point clouds. Therefore, we use the first frame point cloud $P_{1}^{k}$ and the last frame point cloud $P_{N}^{k}$ to initialize the motion axis, performing point cloud registration to derive the transformation matrix between them:
@@ 5.5 Experiment Results
PARIS-scene, PARIS-obj, and DGMarbles* predict motion parameters given motion types and are unable to identify the motion parts from monocular videos. Therefore, we primarily evaluate $AE$ and $PE$. In our dataset, each scene may contain multiple motion parts with different motion types. To enable more intuitive evaluation of the methods’ performance, we categorize all motion parts in the dataset by motion type and object class. The quantitative results are presented in Tab.1. Quantitative comparisons show that, compared to DGMarbles*, our method significantly improves the precision of motion axis parameters. Due to the limited field of view, both PARIS-scene and PARIS-obj fail to capture the objects comprehensively, resulting in poor performance. Fig.4 qualitatively demonstrates the advantages of our method in predicting motion axis.
@@ 6 Conclusion
However, our method still has some limitations. First, similar to other dynamic scene reconstruction techniques, our approach requires individual optimization for each scene, which may limit its application scenarios and scope due to efficiency concerns. Second, our method relies on offline depth estimation and optical flow estimation methods. Although these methods have achieved near-maturity in terms of efficiency and quality, they may perform poorly in certain special scenarios (e.g., highly dynamic scenes), leading to failures in the initial parsing results and thereby restricting the framework’s applicability. We regard these challenges as future research directions to further improve the robustness and practicality of the method.

### v3

@@ Abstract
Accurately analyzing the motion parts and their motion attributes in dynamic environments is crucial for advancing key areas such as embodied intelligence. Addressing the limitations of existing methods that rely on dense multi-view images or detailed part-level annotations, we propose an innovative framework that can analyze 3D mobility from monocular videos in a zero-shot manner. This framework can precisely parse motion parts and motion attributes only using a monocular video, completely eliminating the need for annotated training data. Specifically, our method first constructs the scene geometry and roughly analyzes the motion parts and their initial motion attributes combining depth estimation, optical flow analysis and point cloud registration method, then employs 2D Gaussian splatting for scene representation. Building on this, we introduce an end-to-end dynamic scene optimization algorithm specifically designed for articulated objects, refining the initial analysis results to ensure the system can handle ‘rotation’, ‘translation’, and even complex movements (‘rotation+translation’), demonstrating high flexibility and versatility. To validate the robustness and wide applicability of our method, we created a comprehensive dataset comprising both simulated and real-world scenarios. Experimental results show that our framework can effectively analyze articulated object motions in an annotation-free manner, showcasing its significant potential in future embodied intelligence applications. The project page is at: https://monomobility.github.io/MonoMobility.
@@ 2 Related Work
3D Gaussian-Based Dynamic Scene Reconstruction. 3D Gaussian-based approaches [46, 37, 42, 7, 8] achieve state-of-the-art performance in 3D/4D reconstruction. 4D Gaussian Splatting [46] predicts Gaussian attribute changes over time using a six-plane dynamic module. Subsequent works [37, 42] explore single-view Gaussian dynamic reconstruction. DGMarbles [37] and Shape of Motion [42] leverage advanced data-driven techniques to obtain geometric and motion priors. Unlike these works, which treat each Gaussian independently, our method unifies motion part Gaussians to capture realistic motion patterns.
@@ 4.3 Initialization
In Section 4.2, we have obtained the point cloud sequence $\{P_{i}^{k}\}_{i=1}^{N}$ for each motion part Gaussians $GS^{k}$. However, due to errors in geometric and spatiotemporal alignment of depth estimation results, only when the movement is significant, the motion pattern can be roughly reflected in the transformation between two frame point clouds. Therefore, we use the two point clouds from $\{P_{i}^{k}\}_{i=1}^{N}$ that exhibit the largest motion magnitude to initialize the motion axis, performing point cloud registration to derive the transformation matrix between them:
@@ 5.5 Experiment Results
PARIS-scene, PARIS-obj, and DGMarbles* predict motion parameters given motion types and are unable to identify the motion parts from monocular videos. Therefore, we primarily evaluate $AE$ and $PE$. In our dataset, each scene may contain multiple motion parts with different motion types. To enable more intuitive evaluation of the methods’ performance, we categorize all motion parts in the dataset by motion type and object class. The quantitative results are presented in Tab.1. Quantitative comparisons show that, compared to DGMarbles*, our method significantly improves the precision of motion axis parameters. Due to the limited field of view, both PARIS-scene and PARIS-obj fail to capture the objects comprehensively, resulting in poor performance. Fig.4 qualitatively demonstrates the advantages of our method in predicting motion axis. More experiment results are provided in the supplementary material.
@@ 6 Conclusion
However, our method still has some limitations. First, similar to other dynamic scene reconstruction techniques, our approach requires individual optimization for each scene, which may limit its application scenarios and scope due to efficiency concerns. Second, our method relies on offline depth estimation and optical flow estimation methods. Although these methods have achieved near-maturity in terms of efficiency and quality, they may perform poorly in certain special scenarios, leading to failures in the initial parsing results and thereby restricting the framework’s applicability. More detailed analysis can be found in the supplementary material. We regard these challenges as future research directions to further improve the robustness and practicality of the method.
7 Acknowledge
This work was supported in part by the NSFC (62325211, 62132021), the Major Program of Xiangjiang Laboratory (23XJ01009), the Chongqing Natural Science Foundation (CSTB2024NSCQ-MSX1026), the Fundamental Research Funds for the Central Universities (No. SWU-KT25012), Key R&D Program of Wuhan (2024060702030143).